# Most discriminative stimuli for functional cell type identification

### 29 Nov 2023 v1

@@ Abstract
Identifying cell types and understanding their functional properties is crucial for unraveling the mechanisms underlying perception and cognition. In the retina, functional types can be identified by carefully selected stimuli, but this requires expert domain knowledge and biases the procedure towards previously known cell types. In the visual cortex, it is still unknown what functional types exist and how to identify them. Thus, for unbiased identification of the functional cell types in retina and visual cortex, new approaches are needed. Here we propose an optimization-based clustering approach using deep predictive models to obtain functional clusters of neurons using Most Discriminative Stimuli (MDS). Our approach alternates between stimulus optimization with cluster reassignment akin to an expectation-maximization algorithm. The algorithm recovers functional clusters in mouse retina, marmoset retina and macaque visual area V4. This demonstrates that our approach can successfully find discriminative stimuli across species, stages of the visual system and recording techniques. The resulting most discriminative stimuli can be used to assign functional cell types fast and on the fly, without the need to train complex predictive models or show a large natural scene dataset, paving the way for experiments that were previously limited by experimental time. Crucially, MDS are interpretable: they visualize the distinctive stimulus patterns that most unambiguously identify a specific type of neuron. We will make our code available online upon publication.
@@ 1 Introduction
Animals perceive the world through an intricate network of neurons in the visual system. These neurons are organized into cell types that can be identified by their distinct responses to visual stimuli. Such functional cell typing has been demonstrated in the mouse retina (Baden et al., 2016), where functional cell types also match in their gene expression and morphology (Goetz et al., 2022). In higher visual areas, it has proven difficult to identify cell types based on light responses, although recent attempts point at a mix of continuous and discrete functional types in mouse primary visual cortex (V1) (Ustyuzhaninov et al., 2020; 2022) and primate cortical area V4 (Willeke et al., 2023).
Functional cell type identification currently often requires domain knowledge to handcraft suitable visual stimuli (Farrow & Masland, 2011; Baden et al., 2016; Franke et al., 2017) and select distinctive response features for classification (Farrow & Masland, 2011). This biases classification towards known cell types and might make it hard to discover unknown types. Also, in many brain areas the functional types are still unknown – and hence no “ground truth” types exist to tailor stimuli towards. Therefore, both the study of well known and less understood visual areas would benefit from a more neutral, data-driven approach that does not require domain knowledge. Achieving this goal has recently come within reach: artificial neural networks have been proposed as “digital twins” of the visual system – these networks receive visual stimuli as inputs and predict how a specific biological neuron would respond to that stimulus with high fidelity (Fig. 1A; Antolík et al. (2016); Batty et al. (2016); McIntosh et al. (2016); Klindt et al. (2017); Sinz et al. (2018); Walker et al. (2019); Lurz et al. (2020); Willeke et al. (2022)). Such digital twins make it possible to analyze the computational properties of visual neurons in silico. For instance, digital twins allow to optimize for a stimulus that maximally drives a single neuron, its Maximally Exciting Inputs (MEIs) (Walker et al., 2019; Bashivan et al., 2019), that visualizes which feature the neuron is maximally tuned to.
Bashivan et al. (2019) attempted to make MEIs more specific by optimizing images that do not only drive one neuron, but also suppress all others. However, this is often not possible in practice, because multiple neurons of the same functional cell type are present, which respond similarly.
@@
If we knew which cells belonged to the same functional type, we could easily extend this approach and find Most Discriminative Stimuli (MDS) at the cell type level – stimuli that drive one cell type while suppressing all others. With the resulting MDS, we can discover the unique visual features that the different cell types preferably process and easily identify them in future experiments with the MDS. As it is unknown which neurons belong to which cell type a priori, we need to cluster cells into types and find the MDS for each type in a joint procedure.
Here, we show how to solve this task in an Expectation-Maximization (EM) style manner: We alternate between assigning neurons to cell type clusters, and optimizing an MDS for each cluster (Fig. 1B). We empirically validated our novel clustering algorithm by benchmarking the resulting clusters against well-established retinal ganglion cell (RGC) types (Baden et al., 2016; Field & Chichilnisky, 2007). We further demonstrate that our method found meaningful functional clusters when little domain knowledge is available, as in cortical area V4. Our clustering worked across species (mouse, marmoset, macaque) and recording techniques (two-photon, electrophysiology), and can identify cell types up to $20\%$ faster than traditional methods.
@@ 2 Related work
Cell type clustering. Cell types form the building blocks of neural circuits across the brain (Sanes & Masland, 2015). To understand how these cell types perform computations in the visual system, a functional classification based on light responses may be most relevant (Vlasits et al., 2019). In the mouse retina, functional cell types have been studied extensively (Farrow & Masland, 2011; Baden et al., 2016; Franke et al., 2017), especially for retinal ganglion cells: RGCs can be hierarchically divided by response polarity into ON, OFF and ON-OFF types, and then further by response dynamics into transient and sustained types (Farrow & Masland, 2011). Using more complex stimuli, Baden et al. (2016) identified a finer cluster structure of at least 32 functional RGC types many of which were later matched to morphological and genetic types (Bae et al., 2018; Tran et al., 2019; Goetz et al., 2022; Huang et al., 2022). As the Baden et al. (2016) functional clustering for RGCs is the most comprehensive available to date, we focus on their cluster labels in the retina experiments presented here. In cortex, functional cell types have been harder to define. For instance, the fingerprinting stimuli used in the retina rely mainly on temporal characteristics of full-field brightness changes or known properties such as direction selectivity that can be probed with a few moving bars. How to find appropriate stimuli for the diverse spatial selectivity patterns observed in visual cortex Walker et al. (2019); Ustyuzhaninov et al. (2022); Tong et al. (2023) is less clear. Ustyuzhaninov et al. (2020; 2022) modeled light responses to natural images in mouse primary visual cortex (V1) with a convolutional neural network (CNN), where each neuron’s response is modeled by a linear readout from a shared feature space. They then used the linear readout weights as a vector representation of the neuron’s function. This approach captures the full selectivity of a neuron, but it requires showing natural scenes and training a predictive model to infer a neuron’s cell type, whereas our approach can identify a neuron’s cell type in new experiments quickly and without model training.
@@
Image optimization. Optimizing input images has been used to maximize the response of units in artificial neural networks for object classification (Erhan et al., 2009; Zeiler & Fergus, 2014; Olah et al., 2017) or to distinguish between different classification models (Golan et al., 2020). Building on these ideas, Deep Neural Network (DNN) models of neural activity have been used to generate maximally exciting inputs (MEIs) for biological neurons. These MEIs can drive single neurons in vivo (Bashivan et al., 2019; Walker et al., 2019; Franke et al., 2022; Hoefling et al., 2022; Willeke et al., 2023; Pierzchlewicz et al., 2023), even when using multiple diverse MEIs (Ding et al., 2023), inputs that inhibit a single neuron (Fu et al., 2023), or inputs that drive a single neuron but suppress others (Bashivan et al., 2019). While MEIs were recently extended to drive a predefined group of neurons (Ustyuzhaninov et al., 2022), there is no technique that at the same time suppresses the activity of all other neurons. Here, we provide this missing technique and extend it into a full-fledged functional clustering algorithm.
@@ 3 Most discriminative stimulus clustering algorithm
Digital twins are deep predictive models, trained to predict neural activity based on the stimuli shown. They allow us to search the stimulus space for most discriminative stimuli (MDS) that optimally separate functional groups of neurons in their activity (Fig. 1A). Here, we simultaneously generated MDS and clustered neurons in an EM-like fashion (see Fig. 1B and Algorithm 1 for summary). We started by randomly assigning neurons into a small number of initial “seed” clusters. Then we alternated between (M) optimizing an MDS for each cluster that exclusively activated neurons in that cluster, while avoiding to activate all other clusters, and (E) re-assigning each neuron to the cluster whose MDS drove the neuron the most (Fig. 1B). Periodically, we split candidate clusters into sub-clusters that we kept only if this improved the objective (see below).
@@
where the temperature parameter $\tau$ determines how strongly we penalize large responses in other clusters (we set $\tau=1.6$ which is the best within a broad optimum of the final objective, see Fig. 14). This yielded one MDS per cluster. In the E step, we re-assigned neurons to the cluster associated with the MDS to which they responded most strongly. We then alternated between E- and M-step until cluster assignments did not change anymore.
@@ 4 Data and implementation details
We assigned neurons to Baden et al. (2016) functional types using a classifier that predicts the cell type based on soma size and the light response to two synthetic fingerprinting stimuli, a “chirp” and a moving bar (Qiu et al., 2023). We removed neurons with $<25\%$ assignment confidence and all functional types containing $<50$ modeled neurons, as they might have been underrepresented in digital twin training. Further, displaced amacrine cells, which are usually part of this dataset, were removed as well. This led to 2,448 RGCs of 17 types. We split cells into a training set (80%) and report results on the held-out test set. Splits retained relative functional group sizes. These type labels are no perfect ground truth, as they are subject to potential misassignments by the classifier.
@@
Stimulus optimization. We synthesized MDS with the same parameters that Hoefling et al. (2022) used to optimize MEIs and summarize the key components here. We initialized the stimulus, a 50 frames ×$\times\,(18\times 16)$ pixels ×$\times 2$ channels video-snippet, with Gaussian white noise. We fed this stimulus to the digital twin predicting a 20 frames response trace for each cell. The average predicted activity over the last 10 frames was the input to the objective function (Eq. 1). We optimized the stimulus by gradient ascent for ten steps with a learning rate of 100 (the optimal learning rate within the range $[10^{-3},10^{3}]$). During algorithm development we verified that sufficient optimization steps were performed by inspecting the objective values during optimization. To ensure that the MDS stays within the pixel intensity range used when training the digital twin (Hoefling et al., 2022), we jointly renormalized both channels’ pixel values of each frame to an L2 norm of 30 after each optimization step. Additionally, we clipped the pixel values to the minimum (UV: $-0.913$, green: $-0.654$) and maximum (UV and green: $6.269$) light intensity of the stimulus projector that displayed the videos, where zero corresponds to the mean luminance across the stimuli. This avoids that single pixels can exceed the display range of the projector and is a typical procedure in stimulus optimization settings (e.g. Walker et al. (2019); Hoefling et al. (2022); Willeke et al. (2023)).
@@
Simulation of experiment. To estimate how time-efficient MDS are in determining a functional type, we did an in silico experiment with simulated observational noise. We modeled the trial-to-trial variability for each neuron by a gamma distribution whose mean we set to the digital twin’s response prediction for that neuron. We set the variance to be proportional to the mean and estimated the constant of this proportionality for each neuron. For that, we used the real responses on a subset of the stimuli for which responses to three repeated stimulus presentations were available. Then we presented the MDS to the digital twin, generated 16 “trials” of activity and computed the neural responses by averaging the activity over the last ten frames (as in the stimulus optimization procedure). Model neurons were then assigned to the cluster whose MDS activated them the most.
@@
Marmoset retinal ganglion cells. We additionally evaluated our approach on a dataset of 167 marmoset RGCs obtained from Multi-Electrode Array (MEA) recordings of spiking activity in one marmoset retina while presenting a naturalistic gray-scale movie. For this dataset, no reference taxonomy of functional types is available. We used a digital twin whose architecture generally followed the mouse retina digital twin (Hoefling et al., 2022), but was optimized for this dataset (4 CNN layers, larger 21 pixel spatial filters in the first layer). For digital twin details and MDS optimization parameters, see Sec. A.3.
@@ 5 Results
Figure 2: Most discriminative stimuli clustering on mouse RGCs. Results on held-out test neurons and empirical comparison to Baden et al. (2016) types shown. A. Cluster averaged digital twin response to the optimized MDS. Elements within a column normalized to highest value. B. MDS video snippets per cluster decomposed into time, UV- and green-channel stimulus components by singular value decomposition. For display, spatial components were re-scaled to a symmetric color scale between -1 (blue) and 1 (red). C. Summary of the functional characteristics for each cluster (UV-pref: UV-preference, T-freq. pref.: Temporal frequency preference; Surr.: Surround properties; RF: Receptive Field). D. Confusion matrix between MDS clusters and Baden et al. (2016) types (predicted by a classifier for the mouse RGC dataset; Qiu et al. (2023)). Annotations in same color belong to the same hierarchical functional type. Elements within rows (across Baden et al. (2016) types) were normalized to sum to 1, annotated numbers display the number of neurons in percent. E. Confusion matrix of the classifier (Qiu et al., 2023) used to predict Baden et al. (2016) types for our dataset only having access to the same features as MDS clustering between Baden et al. (2016) type labels and predicted type labels on held out test data. Types confused by the classifier are merged into MDS clusters. Rows normalized (across Baden et al. (2016) types) to sum to 1, annotations displaying the number of neurons in percent.
@@
Mouse retinal ganglion cells. We started by empirically validating our novel clustering algorithm on the mouse RGC dataset that was labeled with 17 well-established Baden et al. (2016) functional types. Our MDS clustering algorithm found seven functional clusters after discarding small clusters with less than ten cells. Clusters were reasonably well separated in terms of their mean response (Fig. 2A) to the MDS (Fig. 2B). While our MDS clustering yielded fewer clusters than that by Baden et al. (2016), the clusters we found comprised the well known functional cluster hierarchy of the retina (e.g., Baden et al. (2016); Farrow & Masland (2011)), namely OFF, ON-OFF, fast ON, and slow ON types, and one color-opponent type that was recently highlighted as having very distinct selectivity (Hoefling et al., 2022). Clusters differed in their preference for temporal frequencies, receptive field center sizes, strength of UV preference, and surround strength and structure (Fig. 2C).
@@
Next, we compared the 17 Baden et al. (2016) types to our MDS clusters (Fig. 2D). We found that groups of Baden et al. (2016) types with similar functional properties matched well with our MDS clusters (color-coded in the labels in Fig. 2D based on clustering hierarchy in Baden et al. (2016)). Off-diagonal entries in Fig. 2A,D show that not all fine types could be perfectly separated by MDS. However, the block-diagonal structure (black boxes in Fig. 2D) indicates that cell types are only confused within groups of very similar function. All OFF and ON-OFF types (orange labels in Fig. 2D) mapped almost exclusively to the MDS clusters 2 and 3, with a tendency of OFF cells mapping to cluster 2, and ON-OFF cells to cluster 3. From the groups of Baden et al. (2016) types, fast ON types (“ON loc tr OS”, “ON trans”, “ON high freq”; green in Fig. 2D) showed the clearest correspondence to a single MDS cluster, 38%, 93%, and 93% of cells of its three types mapping to MDS cluster 4, respectively. Slow ON types (blue in Fig. 2D) mapped mostly to clusters 5 and 6, with a preference for one of the two in most but not all types. In this group, one type (“ON DS sustained”) even mapped exclusively to a single MDS cluster. Despite its name, the “OFF suppressed 1” type also mapped to the ON MDS cluster 6, likely because this type has an ON response for some local stimuli (e.g., see moving bar response in Baden et al. (2016)). The biggest exception of the slow ON types was “ON contrast suppressed”, which did not map strongly to either cluster 5 or 6, but instead mapped to the color opponent MDS cluster 7 (see Fig. 2D). Interestingly, MDS cluster 7 was also the cluster with the strongest 1-to-1 correspondence with a single Baden et al. (2016) type. The only other MDS cluster consisting of mostly one type was MDS cluster 1, which was dominated by the “OFF suppressed 2” type but comprised also a number of other types, including both OFF and ON cells. Similar to “OFF suppressed 2”, cells of “OFF suppressed 1” can be found in most clusters. This is expected, given that both types likely comprise sub-types with distinct functional properties (see coverage factor in Baden et al. (2016)).
@@
We further analyzed why most Baden et al. (2016) types mapped to two or more MDS clusters and why all MDS clusters comprised more than one Baden et al. (2016) type. One hint was that direction- and orientation selective Baden et al. (2016) types (DS and OS in Fig. 2D) were mixed into all three major type groups (colored in Fig. 2D), suggesting the clustering does not capture direction selectivity. Consistent with that idea, we found that none of the MDS exhibited significant lateral motion. This is likely because the used digital twin is limited in capturing direction selectivity by both architecture (space-time-separable convolution) and training data (sampling bias in optical flow; Hoefling et al. (2022)). We note that this is a limitation of the predictive model, not of our procedure. Also, the original Baden et al. (2016) types were constructed with additional information on direction selectivity of each cell that our MDS clustering did not have access to. The MDS clustering can therefore not distinguish cells based on their direction selectivity, and instead merged types with and without direction selectivity when they have otherwise similar functional properties. To confirm this, we re-created cell type labels for a fairer comparison using a reduced classifier. This classifier re-labeled cells using only information the MDS clustering also had access to, namely functional responses to “chirp” stimuli, but not direction selectivity or soma size. We found that the reduced classifier performed substantially worse (accuracy on all types dropped from 79% to 49%) on test data. Furthermore, we found that this reduced classifier confused types that were merged by MDS clustering into single clusters (Fig. 2E). For instance, the reduced classifier confused OFF and ON-OFF types that were mapped to MDS clusters 2 and 3 (Fig. 2D,E). The reduced classifier also confused the fast ON types which were mapped mostly to the single MDS cluster 4, and some of the slow ON types mapping to MDS clusters 5 and 6. This indicates that without access to direction selectivity or morphological features (e.g., soma size) the finer Baden et al. (2016) types cannot be reliably distinguished. This also explains why our MDS clustering finds fewer clusters.
@@
Robustness. To verify that our clustering algorithm returned consistent results across runs, we repeated MDS clustering with another random initialization of the stimulus and initial neuron assignments into 30 instead of 5 clusters. Again the algorithm converged to seven clusters, which exhibited strikingly similar MDS as the run initialized with five clusters (Appendix Fig. 6). The similarity of clusters assignments across eleven comparison runs varying in initial number of clusters, initial neuron assignment, and MDS initialization was also high (median ARI: 0.67), suggesting MDS clustering is robust across initializations. Further, running MDS clustering for a different digital twin architecture yielded similar results, suggesting robustness to the specific digital twin model used (see Sec. A.6 and Fig. 13 for details).

### 14 Mar 2024 v2

# Most discriminative stimuli for functional cell type clustering

@@ Abstract
Identifying cell types and understanding their functional properties is crucial for unraveling the mechanisms underlying perception and cognition. In the retina, functional types can be identified by carefully selected stimuli, but this requires expert domain knowledge and biases the procedure towards previously known cell types. In the visual cortex, it is still unknown what functional types exist and how to identify them. Thus, for unbiased identification of the functional cell types in retina and visual cortex, new approaches are needed. Here we propose an optimization-based clustering approach using deep predictive models to obtain functional clusters of neurons using Most Discriminative Stimuli (MDS). Our approach alternates between stimulus optimization with cluster reassignment akin to an expectation-maximization algorithm. The algorithm recovers functional clusters in mouse retina, marmoset retina and macaque visual area V4. This demonstrates that our approach can successfully find discriminative stimuli across species, stages of the visual system and recording techniques. The resulting most discriminative stimuli can be used to assign functional cell types fast and on the fly, without the need to train complex predictive models or show a large natural scene dataset, paving the way for experiments that were previously limited by experimental time. Crucially, MDS are interpretable: they visualize the distinctive stimulus patterns that most unambiguously identify a specific type of neuron.
@@ 1 Introduction
Animals perceive the world through an intricate network of neurons in the visual system. These neurons are organized into cell types that can be identified by their distinct responses to visual stimuli. Such functional cell typing has been demonstrated in the mouse retina (Baden et al., 2016), where functional cell types also match in their gene expression and morphology (Goetz et al., 2022). In higher visual areas, it has proven difficult to identify cell types based on their response patterns, although recent attempts point at a mix of continuous and discrete functional types in mouse primary visual cortex (V1) (Ustyuzhaninov et al., 2020; 2022) and primate cortical area V4 (Willeke et al., 2023).
Functional cell type identification currently often requires domain knowledge to handcraft suitable visual finger-printing stimuli (Farrow & Masland, 2011; Baden et al., 2016; Franke et al., 2017) and to select distinctive response features for classification (Farrow & Masland, 2011). This biases classification towards known cell types and might make it hard to discover new types. Also, in many brain areas the functional types are still unknown – hence no “ground truth” exists to tailor stimuli towards. Therefore, the study of both well known and less understood visual areas would benefit from a more neutral, data-driven approach that does not require domain knowledge. One idea would be to directly cluster cell’s responses to unbiased naturalistic stimuli. However, the receptive field locations vary within a simultaneously recorded population of neurons, which means that each cell would “see” a different part of the stimulus. Hence, even cells with identical input-output function will respond differently to the same stimulus, preventing direct clustering of their responses.
Fortunately, resolving this issue has recently come within reach: artificial neural networks have been proposed as “digital twins” of the visual system – these networks receive visual stimuli as inputs and predict how a specific biological neuron would respond to that stimulus with high fidelity (Fig. 1A; Antolík et al. (2016); Batty et al. (2016); McIntosh et al. (2016); Klindt et al. (2017); Sinz et al. (2018); Walker et al. (2019); Lurz et al. (2020); Burg et al. (2021); Willeke et al. (2022); Ustyuzhaninov et al. (2022)). Such digital twins make it possible to analyze the computational properties of visual neurons in silico. For instance, digital twins allow to center a stimulus on each individual neuron’s receptive field (Ustyuzhaninov et al., 2022) and to optimize it to maximally drive a single neuron, yielding its Maximally Exciting Input (MEI), that visualizes the neuron’s preferred feature (Walker et al., 2019; Bashivan et al., 2019). Bashivan et al. (2019) attempted to make MEIs more specific by optimizing images that do not only drive one neuron, but also suppress all others. However, this is often not possible in practice, because multiple neurons of the same functional cell type are present, which respond similarly.
@@
If we knew which cells belong to the same functional type, we could easily extend this approach and find Most Discriminative Stimuli (MDS) at the cell type level – stimuli that drive one cell type while suppressing all others when presented centered on each cell’s receptive field. With the resulting MDS, we can discover the unique visual features that the different cell types preferably process, rendering MDS to be efficient finger-printing stimuli for easy cell type identification in future experiments. As it is unknown which neurons belong to which cell type a priori, we need to cluster cells into types and find the MDS for each type in a joint procedure.
Here, we tackle this task in an Expectation-Maximization (EM) style manner: We alternate between assigning neurons to cell type clusters, and optimizing a MDS for each cluster (Fig. 1B). We empirically validate our novel clustering algorithm by benchmarking the resulting clusters against well-established retinal ganglion cell (RGC) types (Baden et al., 2016; Field & Chichilnisky, 2007). We further demonstrate that our method found meaningful functional clusters when little domain knowledge is available, as in cortical area V4. Our clustering works across species (mouse, marmoset, macaque) and recording techniques (two-photon imaging, electrophysiology), and leveraging MDS as finger-printing stimuli can identify cell types $20\%$ faster than traditional methods.
@@ 2 Related work
Cell type clustering. Cell types form the building blocks of neural circuits across the brain (Sanes & Masland, 2015). To understand how these cell types perform computations in the visual system, a functional classification based on light responses may be most relevant (Vlasits et al., 2019). In the mouse retina, functional cell types have been studied extensively (Farrow & Masland, 2011; Baden et al., 2016; Franke et al., 2017), especially for retinal ganglion cells: RGCs can be hierarchically divided by response polarity into ON, OFF and ON-OFF types, and then further by response dynamics into transient and sustained types (Farrow & Masland, 2011). Using more complex stimuli, Baden et al. (2016) identified a finer cluster structure of at least 32 functional RGC types, many of which were later matched to morphological and genetic types (Bae et al., 2018; Tran et al., 2019; Goetz et al., 2022; Huang et al., 2022). As the Baden et al. (2016) functional clustering for RGCs is the most comprehensive available to date, we focus on their cluster labels in the retina experiments presented here. In cortex, functional cell types have been harder to define. For instance, the fingerprinting stimuli used in the retina rely mainly on temporal characteristics of full-field brightness changes or known properties such as direction selectivity probed with moving bars. How to find appropriate stimuli for the diverse spatial selectivity patterns observed in visual cortex (Walker et al., 2019; Ustyuzhaninov et al., 2022; Tong et al., 2023) is less clear. Ustyuzhaninov et al. (2020; 2022) modeled responses to natural images in mouse primary visual cortex (V1) with a convolutional neural network (CNN), where each neuron’s response is modeled by a linear readout from a shared feature space. They then used the readout weights as a vector representation of the neuron’s function. This approach captures the full selectivity of a neuron, but it requires showing natural scenes and training a predictive model to infer a neuron’s cell type, whereas our approach can identify a neuron’s cell type in new experiments quickly and without model training.
@@
Image optimization. Optimizing input images has been used to maximize the response of units in artificial neural networks for object classification (Erhan et al., 2009; Zeiler & Fergus, 2014; Olah et al., 2017) and to distinguish between different classification models (Golan et al., 2020). Building on these ideas, Deep Neural Network (DNN) models of neural activity were used to generate maximally exciting inputs (MEIs) for biological neurons. These MEIs can drive single neurons in vivo (Bashivan et al., 2019; Walker et al., 2019; Franke et al., 2022; Hoefling et al., 2022; Willeke et al., 2023; Pierzchlewicz et al., 2023), even when using multiple diverse MEIs (Ding et al., 2023); inhibit a single neuron (Fu et al., 2023), or drive a single neuron but suppress others (Bashivan et al., 2019). While MEIs were recently extended to drive a predefined group of neurons (Ustyuzhaninov et al., 2022), there is no technique that at the same time suppresses the activity of all other neurons. Here, we provide this missing technique and extend it into a full-fledged clustering algorithm.
@@ 3 Most discriminative stimulus clustering algorithm
Digital twins are models trained to predict neural activity based on shown stimuli. They allow us to search the stimulus space for most discriminative stimuli (MDS) that optimally separate functional groups of neurons in their activity (Fig. 1A). Here, we simultaneously generated MDS and clustered neurons in an EM-like fashion (Fig. 1B and Algorithm 1). We started by randomly assigning neurons into a small number of initial “seed” clusters. Then we alternated between (M) optimizing an MDS for each cluster exclusively activating that cluster’s neurons, while avoiding to activate other clusters, and (E) re-assigning neurons to the cluster whose MDS drove them the most (Fig. 1B). Periodically, we split candidate clusters into sub-clusters that we kept if this improved the objective (see below).
@@
where the temperature parameter $\tau$ determines how strongly we penalize large responses in other clusters (we set $\tau=1.6$ which is the best within a broad optimum of the final objective, see Fig. 14). This yielded one MDS per cluster. In the E step, we re-assigned neurons to the cluster associated with the MDS to which they responded most strongly. We then alternated between the E- and M-step until cluster assignments did not change anymore.
@@ 4 Data and implementation details
We assigned neurons to Baden et al. (2016) functional types using a classifier that predicts the cell type based on soma size and the light response to two synthetic fingerprinting stimuli, a “chirp” and a moving bar (Qiu et al., 2023). We removed neurons with less than $25\%$ assignment confidence and all functional types containing less than $50$ modeled neurons, as they might have been underrepresented in digital twin training. Further, displaced amacrine cells, which are usually part of this dataset, were removed as well. This led to 2,448 RGCs of 17 types. To ensure MDS are not overfit to the cells they were optimized on and that they also work in future experiments, we computed MDS on a 80% training split of the cells and report results on the remaining held-out test cells. Splits retained relative functional group sizes. These type labels are no perfect ground truth, as they are subject to potential misassignments by the classifier.
@@
Stimulus optimization. We synthesized MDS with the same parameters that Hoefling et al. (2022) used to optimize MEIs and summarize the key components here. We initialized the stimulus, a 50 frames ×$\times\,(18\times 16)$ pixels ×$\times 2$ channels video-snippet, with Gaussian white noise. We fed this stimulus to the digital twin predicting a 20 frames response trace for each cell. The average predicted activity over the last ten frames was the input into the objective function (Eq. 1). We optimized the stimulus by gradient ascent for ten steps with a learning rate of 100 (the optimal learning rate within the range $[10^{-3},10^{3}]$). During algorithm development we verified that sufficient optimization steps were performed by inspecting the objective values during optimization. To ensure that the MDS stayed within the pixel intensity range used when training the digital twin (Hoefling et al., 2022), we jointly renormalized both channels’ pixel values of each frame to an $L_{2}$-norm of 30 after each optimization step. Additionally, we clipped the pixel values to the minimum (UV: $-0.913$, green: $-0.654$) and maximum (UV and green: $6.269$) light intensity of the stimulus projector that displayed the videos, where zero corresponds to the mean luminance across the stimuli. This avoids that single pixels can exceed the display range of the projector and is a typical procedure in stimulus optimization settings (e.g. Walker et al. (2019); Hoefling et al. (2022); Willeke et al. (2023)).
@@
Simulation of experiment. To estimate how time-efficient MDS are in determining a functional type, we did an in silico experiment with simulated observational noise. We modeled the trial-to-trial variability for each neuron by a gamma distribution whose mean we set to the digital twin’s MDS response prediction for that neuron. We set the variance to be proportional to the mean and estimated the proportionality constant for each neuron from a subset of the stimuli in the dataset for which responses to three repeated stimulus presentations were provided. Then we sampled from the gamma distribution to generate 16 “trials” of activity and computed the neural responses by averaging the activity over the last ten frames (as in the stimulus optimization procedure). Model neurons were then assigned to the cluster whose MDS activated them the most.
@@
Marmoset retinal ganglion cells. We additionally evaluated our approach on 167 marmoset RGCs obtained from Multi-Electrode Array (MEA) recordings of spiking activity in one retina while presenting a naturalistic gray-scale movie. For this data, no fine-grained reference taxonomy of functional types exists. We optimized the mouse digital twin (Hoefling et al., 2022) for this dataset (four CNN layers, larger 21 pixel spatial filters in the first layer; see Sec. A.3 for details).
@@ 5 Results
[…]
@@
Mouse retinal ganglion cells. We started by empirically validating our novel clustering algorithm on the mouse RGC dataset that was labeled with 17 well-established Baden et al. (2016) functional types. Our MDS clustering algorithm found seven functional clusters after discarding small clusters with less than ten cells. Clusters were reasonably well separated in terms of their mean response (Fig. 2A) to the MDS (Fig. 2B). While our MDS clustering yielded fewer clusters than that by Baden et al. (2016), the clusters we found comprised the well known functional cluster hierarchy of the retina (e.g. Baden et al. (2016); Farrow & Masland (2011)), namely OFF, ON-OFF, fast ON, and slow ON types, and one color-opponent type that was recently highlighted as having very distinct selectivity (Hoefling et al., 2022). Clusters differed in their preference for temporal frequencies, receptive field center sizes, strength of UV preference, and surround strength and structure (Fig. 2C).
@@
Next, we compared the 17 Baden et al. (2016) types to our MDS clusters (Fig. 2D). We found that groups of Baden et al. (2016) types with similar functional properties matched well with our MDS clusters (color-coded labels in Fig. 2D based on clustering hierarchy in Baden et al. (2016)). Off-diagonal entries in Fig. 2A,D show that not all fine types could be perfectly separated by MDS. However, the block-diagonal structure (black boxes in Fig. 2D) indicates that cell types were only confused within groups of very similar function. All OFF and ON-OFF types (orange labels in Fig. 2D) mapped almost exclusively to MDS clusters 2 and 3, with a tendency of OFF cells mapping to cluster 2, and ON-OFF cells to cluster 3. From the groups of Baden et al. (2016) types, fast ON types (“ON loc tr OS”, “ON trans”, “ON high freq”; green in Fig. 2D) showed the clearest correspondence to a single MDS cluster, 38%, 93%, and 93% of cells of its three types mapping to MDS cluster 4, respectively. Slow ON types (blue in Fig. 2D) mapped mostly to clusters 5 and 6, with a preference for one of the two in most but not all types. In this group, one type (“ON DS sustained”) even mapped exclusively to a single MDS cluster. Despite its name, the “OFF suppressed 1” type also mapped to ON MDS cluster 6, likely because this type has an ON response for some local stimuli (e.g. moving bar response in Baden et al. (2016)). The biggest exception of the slow ON types was “ON contrast suppressed”, which did not map strongly to either cluster 5 or 6, but instead mapped to the color opponent MDS cluster 7 (Fig. 2D). Interestingly, this cluster showed the strongest one-to-one correspondence with a single Baden et al. (2016) type. The only other MDS cluster consisting of mostly one type was MDS cluster 1, which was dominated by the “OFF suppressed 2” type, but also comprised other types, including both OFF and ON cells. Similar to “OFF suppressed 2”, cells of “OFF suppressed 1” could be found in most clusters. This is expected, given that both types likely comprise sub-types with distinct functional properties (see coverage factor in Baden et al. (2016)).
@@
We further analyzed why most Baden et al. (2016) types mapped to two or more MDS clusters and why all MDS clusters comprised more than one Baden et al. (2016) type. One hint was that direction- and orientation selective Baden et al. (2016) types (DS and OS in Fig. 2D) were mixed into all three major cluster groups (colored in Fig. 2D), suggesting the clustering did not capture direction selectivity. Consistent with that idea, we found that none of the MDS exhibited significant lateral motion. This is likely because the used digital twin is limited in capturing direction selectivity by both architecture (space-time-separable convolution) and training data (sampling bias in optical flow; Hoefling et al. (2022)). We note that this is a limitation of the predictive model, not of our procedure. Also, the original Baden et al. (2016) types were constructed with additional information on direction selectivity of each cell that our MDS clustering did not have access to. The MDS clustering can therefore not distinguish cells based on their direction selectivity, and instead merged types with and without direction selectivity when they had otherwise similar functional properties. To confirm this, we re-created cell type labels for a fairer comparison using a reduced classifier. This classifier re-labeled cells using only information the MDS clustering also had access to, namely functional responses to “chirp” stimuli, but not direction selectivity or soma size. We found that the reduced classifier performed substantially worse (accuracy dropped from 79% to 49% on test data) and also confused types that were merged by MDS clustering into single clusters (Fig. 2E). For instance, the reduced classifier confused OFF and ON-OFF types that were mapped to MDS clusters 2 and 3, the fast ON types which were mapped mostly to the single MDS cluster 4, and some of the slow ON types mapping to MDS clusters 5 and 6 (Fig. 2D,E). This indicates that without access to direction selectivity or morphological features (e.g. soma size) the finer Baden et al. (2016) types cannot be reliably distinguished. This also explains why our MDS clustering finds fewer clusters.
@@
Robustness. To verify that our clustering algorithm returned consistent results across runs, we repeated MDS clustering with another random initialization of the stimulus and initial neuron assignments into 30 instead of five clusters. Again the algorithm converged to seven clusters, which exhibited strikingly similar MDS as the run initialized with five clusters (Appendix Fig. 6). The similarity of cluster assignments across eleven comparison runs varying in initial number of clusters, initial neuron assignment, and MDS initialization was also high (median ARI: 0.67), suggesting MDS clustering is robust across initializations. Further, running MDS clustering for a different digital twin architecture yielded similar results, suggesting robustness to the specific digital twin model used (see Sec. A.6 and Fig. 13 for details).